# Bytes to Schlep? Use a FEP: Hiding Protocol Metadata with Fully Encrypted Protocols

### v1

@@ H.1. Experimental Setup
Unless otherwise stated, experiments were performed in a Pop! OS 22.04 Linux environment running Linux Kernel 6.6.10, and messages were sent locally, via the IPv4 loopback interface. We wrote custom software to run our experiments. In the datastream setting, each protocol consists of two applications, which we refer to as a Server Proxy and Client Proxy. We wrote a client, server, and an intermediate meddler-in-the-middle (MiTM) proxy to modify bytes in transit between the Client Proxy and Server Proxy. The MiTM proxy can be configured with the direction of the message to modify (towards the server, which we referred to as ‘out’, ‘in’ otherwise), the message to modify (specifying the ordinal $n$ to modify the $n$th message), and the byte within the message to change. By default, we modified the target byte by replacing it with its XOR by the value 1. We introduce two additional transparent MiTM proxies between the server and client themselves and their proxy applications which relayed all connections and messages unchanged, to facilitate easy analysis of packet captures (thus, every connection in the experiment is connected to at least one static, bound port, making it trivial to later determine the sender and receiver of any message or fragment by its source and destination ports alone). Our experiments initialized all MiTM and proxy servers in reverse order (beginning with Server and ending with the MiTM between the Client and Client Proxy). Then we ran the client application, which sent a sequence of ten messages, each with a user-determined quantity of a distinct plaintext byte. The server responded to each client message with by sending a message with the same repeated byte it received in a different quantity. By default to the client sent 600 bytes in each message, and the server sent 500. We chose these values since they were able to be encapsulated in a single message or ciphertext fragment for all of our analyzed protocols, were unlikely to be fragmented or merged by the operating system due to their size, and we chose the client and server length to be distinct to make packet captures easy to read. During experiments to validate minimum lengths, the client and server both sent instead messages of length 1. We ran TShark during all experiments listening on the loopback IPv4 interface writing to a file. The packet captures included the local timestamp for each frame, source and destination ports, message lengths, and the TCP message type (with specific fields for TCP FIN and TCP RST). After the experiment concluded, we wrote software to analyze the packet capture and experiment log, extract the packets within the valid time window for each individual experiment, and identify the sender and receiver of the first TCP FIN or RST frame by the destination and source port. Since many protocols include some handshake or initial state which transitions later to a data transport phase, we ran our experiments both modifying the initial few messages (n=1,2,3) and also messages later in the stream (n=5,10). Once we identified a close behavior, we ran the experiment multiple times, modifying different bytes in the fragment to confirm the result before indicating the close behavior on Table 1.
@@ H.2. Protocols
Shadowsocks-libev. Shadowsocks-libev (Windy et al., 2020) is an implementation of Shadowsocks (sha, 2022) available in the official Debian repository. It is written in C and designed to be lightweight. Shadowsocks is an encrypted SOCKS5 proxy that requires a pre-shared key. The client begins the protocol by transmitting an IV, after which messages are AEAD encrypted, with each ciphertext block prepended by a two-byte length block, which is itself also AEAD encrypted. The Shadowsocks-libev source code explicitly includes logic in the design of the server to handle close behavior by way of a special state called STAGE_STOP (see e.g. Line 971 of src/server.c) which is entered after receiving a message that fails to decrypt, during which all messages are ignored and the connection is held open until terminated by the client. We deployed Shadowsocks-libev version 3.3.5 using the ChaChaPoly1305 cipher suite as recommended in the documentation. For channel closures, we observed that in the outgoing direction, the server kept the connection open without terminating the connection throughout the length of the experiment, but in the incoming direction (i.e. messages directed at the client), the client immediately terminates the TCP connection upon receiving a modified message (whether the modification occurs in the payload or tag of either block in a given message). Shadowsocks does produce error messages when messages fail to authenticate, and so its behavior in the outgoing direction does not technically satisfy FEP-CCFA, but with minor protocol modifications or an application that simply ignores errors, the definition is plausibly satisfied. Shadowsocks does not pad its messages. Since the authentication tag for this stream cipher has length 16, we expected the minimum length of an output fragment under this protocol to be 35 (two tags, two bytes of length data, and a one byte payload) and indeed observed this behavior in our experiments.
@@
V2Ray Framework. The V2Ray framework (v2ray, 2024) implements a variety of protocols for the purposes of censorship circumvention including two FEPs: VMess (vmess, 2024), and Shadowsocks. Shadowsocks is an alternate implementation of the same protocol described above, while VMess is a complex custom FEP which includes many anti-censorship features. It consists of an early phase, where initial and header messages are sent in both directions, followed by a data transport phase. We assessed the both protocols as implemented in the V2Ray framework version 5.7.0. and configured both protocols using the sample configuration files. The V2Ray framework implements a so-called “drain” approach to frustrate active probing attempts, wherein a random number of bytes are selected for each connection, randomized on a per-user and per-connection basis, and the connection is not dropped until after this number of bytes have been read throughout the duration of the connection. In the course of our experiments on these V2Ray protocols we identified behavior that violates traditional security notions for encrypted protocols: they do not satisfy integrity. We observed that during the data transport phase of each protocol, if a byte within the ciphertext containing data is modified so that the authentication tag fails to validate, the payload is dropped but the connection is not closed, meaning that both protocols continue to parse new messages, so an active network adversary can by modifying appropriate bytes in transit drop certain messages from the stream. We disclosed our results and in particular this vulnerability to the framework maintainers on January 4, 2024 including a minimal working example of the attack. Thus, while both protocols satisfy FEP-CPFA, neither can satisfy FEP-CCFA since they do not satisfy INT-CST.
@@
(lib)InterMAC. (lib)InterMAC is a protocol proposed by Boldyreva et al. (Boldyreva et al., 2012) and implemented by Albrecht et al. (Albrecht et al., 2019) which modifies the SSH Binary Packet Protocol in order to provide security properties associated with ciphertext lengths and fragmentation. The protocol transmits encrypted and authenticated fixed-sized chunks with the chunk size predetermined as a configuration parameter, padding to achieve this length if necessary. In response to our queries, the maintainer updated the libInterMAC repository with a complete implementation of the protocol on September 27, 2023, which is the version we tested. When a chunk fails to authenticate, libInterMAC terminates the connection in either direction as we verified in our experiments. We note that this protocol is not in use, and only encrypts the data transport phase of an SSH connection meaning that existing InterMAC implementation is not properly a FEP, but we considered it valuable to include it in our results as an example of a distinct and reasonable approach to this problem from the others discussed here.
Obfs4. Obfs4 (obfs4, 2024) is a fully encrypted transport protocol for Tor (Dingledine et al., 2004). It is a complex protocol that includes a key exchange and resists active probing attacks. The data transport phase of the protocol includes AEAD encrypted ciphertext blocks prepended by a two byte unauthenticated length field which is encrypted by a stream cipher. We installed Obfs4Proxy version 0.1.3 from the official Debian repository and used ptadapter (pta, 2020) to run the transport as a standalone proxy. When a ciphertext fails to decrypt, the connection is immediately terminated by the receiver. This also occurs if an adversary modifies the length field in transit and the resulting expected ciphertext of an incorrect length fails to decrypt, which we were able to verify in our experiments. Thus while the protocol plausibly satisfies INT-CST (if a length field is modified, the protocol produces only channel closures/errors) it does not satisfy FEP-CCFA, since this close function is not secure (it requires knowledge of the encrypted lengths).
@@
OpenVPN-XOR. OpenVPN (OpenVPN, 2024) is a VPN protocol that can be patched to encrypt all outputs by XOR with the repetition of a pre-shared key (OpenVPN XORPatch, 2018). We analyzed OpenVPN version 2.6.8 as provided from the official Debian repository without the XOR patch applied (and assume that the XOR encryption should not modify message lengths) using the parameters set in the example configurations provided with the software, which deploy the proxy using the AES256-GCM cipher suite and with TLS mode turned on, with other configuration options mostly remaining as default. The obfuscation method in place is trivially identifiable and other work (Xue et al., 2022) has carefully studied its close behavior, which is to terminate the connection after an authentication failure. We explain the structure of OpenVPN messages as described in the official documentation (ope, 2024) under these default settings. OpenVPN messages are either data channel messages or control channel messages. Control channel messages are a minimum of 40 bytes if SHA-1 is used for the protocol’s HMAC (which is the default) and if the 4-byte timestamps intended to combat replay attacks are switched off (in our installation, they are on by default). This means adopting the default settings, control channel messages are 44 bytes. Version 2 data channel messages under these settings include a 26 byte header, including a two-byte length field. We did not observe Version 1 data channel messages in our experiments under the default configuration. OpenVPN typically encapsulates other protocols in the data channel, including their headers, but also directly sends messages of its own, including a fixed 16-byte ping message (explicitly defined in src/openvpn/ping.c Line 42), which was the smallest we observed in our experiments. Thus, the minimum length we identified in the TCP setting was 42.
@@
Obfuscated OpenSSH. Obfuscated OpenSSH (GitHub, 2010) is a patch to OpenSSH wherein key material is transmitted in the clear (or a PSK is used), after which the handshake protocol is fully encrypted (the binary packet protocol for SSH is fully encrypted without any modification). The approach of Obfuscated OpenSSH is used in the Psiphon (Psiphon, 2024) anti-censorship application, designed for Windows and mobile devices. We analyzed Obfuscated OpenSSH as downloaded from the official repository, which is a modification of OpenSSH 5.2_p1 but since the protocol includes complex handshake logic we ran it directly through a MiTM proxy to modify bytes in transit through the loopback interface and analyzed the resulting packet captures with Wireshark. OpenSSH terminates the connection immediately in both directions when a binary packet protocol message fails to authenticate, and we found in our experiments that this is no different after the obfuscation patch is applied. OpenSSH does not pad message lengths though this is permitted in the SSH specification. According to the RFC for the binary packet protocol, ”The minimum size of a packet is 16 (or the cipher block size, whichever is larger) bytes (plus ’mac’)” (Lonvick and Ylonen, 2006). For the initial messages, before the session key is negotiated, the MAC field is empty, meaning we expected the minimum to be 16 bytes. We ran the protocol in obfuscated and non-obfuscated configuration, and identified 16 byte messages appearing as part of the initial handshake in both circumstances. in the non-obfuscated captures, we identified this as a New Keys message, which indicates the switch to the recently negotiated session keys.
@@
kcptun. kcptun (kcptun, 2024) is a custom datastream transport protocol available in the official Debian repository, implemented via UDP datagrams and designed for very noisy networks. The protocol uses error correcting codes and sends many redundant messages to achieve a high probability that the sequence of messages can be reconstructed on the receiver side even in degraded network conditions. kcptun packets are by default encrypted using AES in Cipher Feedback mode, with a Cyclic Redundancy Check applied to the plaintext and prepended before encryption, providing non-cryptographic authentication, meaning that while the messages are fully encrypted, the protocol does not satisfy FEP-CPFA since it does not satisfy INT-CST. When the CRC is invalid, the packet is dropped, and kcptun measures the number of dropped packets and adjusts its internal parameters accordingly, significantly and predictably increasing the rate at which redundant messages are sent in direct proportion to the network packet loss. We wrote distinct software to proxy UDP messages, and did not notice a change in protocol behavior with increased induced packet loss beyond a predictable change in volume of packet transmissions. kcptun does not include padding and uses empty messages as keepalives which are the smallest that arose in our experiments, which are 16 (nonce) + 4 (CRC) + 24 (KCP headers) + 8 (Error Correction metadata) bytes for a total of 52 byte keepalive messages. The message format for kcptun messages is complex: a more thorough description of the headers, encapsulation format, and final message layout is given in the project page for the dependency kcp-go (kcp, 2024).
@@
SWGP-go. SWGP (swg, 2024) is a proxy designed for the Wireguard VPN protocol, intended to transform Wireguard into a datagram FEP. It behaves in two modes: one called “zero overhead”, where each Wireguard packet has its first 16 bytes encrypted via AES-ECB using a pre-shared key, and “paranoid”, which pads and AEAD encrypts each message. We assess the protocol in its paranoid mode. We deployed SWGP Version 1.5.0 configured with the default settings suggested by the maintainers on the project landing page. Since we suspected that running both multiple Wireguard and SWGP instances locally would be cumbersome, we ran our experiments through Docker-Compose (doc, 2024) and constructed three Docker images, a client image (including the client Wireguard instance and the client SWGP application), a router image which passed on messages transparently while recording them via TShark, and a server image. The client image also included a Python program to send an empty, raw IP packet through the interface every 5 seconds. Each experiment ran for 5 minutes and we assessed the resulting packet captures recorded by the router image. We observed in the source code that the paranoid protocol padded each message to a random length between one more than the packet’s minimum length and the MTU, which is set as a configuration option and is 1500 as default. Thus we hypothesized the minimum lengths to be the smallest messages in Wireguard, which are 32 byte keepalives, encapsulated with a 24 byte nonce and a 16 byte authentication tag (the encapsulation format is given in packet/paranoid.go, Line 18) with one required minimum padding byte for a total of 75, which is the smallest we observed in our experiments.

### v2

@@ H.1. Experimental Setup
Unless otherwise stated, experiments were performed in a Pop! OS 22.04 Linux environment running Linux Kernel 6.6.10, and messages were sent locally, via the IPv4 loopback interface. We wrote custom software to run our experiments which we make publicly available at https://github.com/efenske/bytes_to_schlep/. In the datastream setting, each protocol consists of two applications, which we refer to as a Server Proxy and Client Proxy. We wrote a client, server, and an intermediate meddler-in-the-middle (MiTM) proxy to modify bytes in transit between the Client Proxy and Server Proxy. The MiTM proxy can be configured with the direction of the message to modify (towards the server, which we referred to as ‘out’, ‘in’ otherwise), the message to modify (specifying the ordinal $n$ to modify the $n$th message), and the byte within the message to change. By default, we modified the target byte by replacing it with its XOR by the value 1. We introduce two additional transparent MiTM proxies between the server and client themselves and their proxy applications which relayed all connections and messages unchanged, to facilitate easy analysis of packet captures (thus, every connection in the experiment is connected to at least one static, bound port, making it trivial to later determine the sender and receiver of any message or fragment by its source and destination ports alone). Our experiments initialized all MiTM and proxy servers in reverse order (beginning with Server and ending with the MiTM between the Client and Client Proxy). Then we ran the client application, which sent a sequence of ten messages, each with a user-determined quantity of a distinct plaintext byte. The server responded to each client message by sending a message with the same repeated byte it received in a different quantity. By default to the client sent 600 bytes in each message, and the server sent 500. We chose these values since they were able to be encapsulated in a single message or ciphertext fragment for all of our analyzed protocols, were unlikely to be fragmented or merged by the operating system due to their size, and we chose the client and server length to be distinct to make packet captures easy to read. During experiments to validate minimum lengths, the client and server both sent instead messages of length 1. We ran TShark during all experiments listening on the loopback IPv4 interface writing to a file. The packet captures included the local timestamp for each frame, source and destination ports, message lengths, and the TCP message type (with specific fields for TCP FIN and TCP RST). After the experiment concluded, we wrote software to analyze the packet capture and experiment log, extract the packets within the valid time window for each individual experiment, and identify the sender and receiver of the first TCP FIN or RST frame by the destination and source port. Since many protocols include some handshake or initial state which transitions later to a data transport phase, we ran our experiments both modifying the initial few messages (n=1,2,3) and also messages later in the stream (n=5,10). Once we identified a close behavior, we ran the experiment multiple times, modifying different bytes in the fragment to confirm the result before indicating the close behavior on Table 1.
@@ H.2. Protocols
Shadowsocks-libev. Shadowsocks-libev (Windy et al., 2020) is an implementation of Shadowsocks (Shadowsocks., 2022) available in the official Debian repository. It is written in C and designed to be lightweight. Shadowsocks is an encrypted SOCKS5 proxy that requires a pre-shared key. The client begins the protocol by transmitting an IV, after which messages are AEAD encrypted, with each ciphertext block prepended by a two-byte length block, which is itself also AEAD encrypted. The Shadowsocks-libev source code explicitly includes logic in the design of the server to handle close behavior by way of a special state called STAGE_STOP (see e.g. Line 971 of src/server.c) which is entered after receiving a message that fails to decrypt, during which all messages are ignored and the connection is held open until terminated by the client. We deployed Shadowsocks-libev version 3.3.5 using the ChaChaPoly1305 cipher suite as recommended in the documentation. For channel closures, we observed that in the outgoing direction, the server kept the connection open without terminating the connection throughout the length of the experiment, but in the incoming direction (i.e. messages directed at the client), the client immediately terminates the TCP connection upon receiving a modified message (whether the modification occurs in the payload or tag of either block in a given message). Shadowsocks does produce error messages when messages fail to authenticate, and so its behavior in the outgoing direction does not technically satisfy FEP-CCFA, but with minor protocol modifications or an application that simply ignores errors, the definition is plausibly satisfied. Shadowsocks does not pad its messages. Since the authentication tag for this stream cipher has length 16, we expected the minimum length of an output fragment under this protocol to be 35 (two tags, two bytes of length data, and a one byte payload) and indeed observed this behavior in our experiments.
@@
V2Ray Framework. The V2Ray framework (Project V., 2024) implements a variety of protocols for the purposes of censorship circumvention including two FEPs: VMess (VMess., 2024), and Shadowsocks. Shadowsocks is an alternate implementation of the same protocol described above, while VMess is a complex custom FEP which includes many anti-censorship features. It consists of an early phase, where initial and header messages are sent in both directions, followed by a data transport phase. We assessed both protocols as implemented in the V2Ray framework version 5.7.0. and configured both protocols using the sample configuration files. The V2Ray framework implements a so-called “drain” approach to frustrate active probing attempts, wherein a random number of bytes are selected for each connection, randomized on a per-user and per-connection basis, and the connection is not dropped until after this number of bytes have been read throughout the duration of the connection. In the course of our experiments on these V2Ray protocols we identified behavior that violates traditional security notions for encrypted protocols: they do not satisfy integrity. We observed that during the data transport phase of each protocol, if a byte within the ciphertext containing data is modified so that the authentication tag fails to validate, the payload is dropped but the connection is not closed, meaning that both protocols continue to parse new messages, so an active network adversary can by modifying appropriate bytes in transit drop certain messages from the stream. We disclosed our results and in particular this vulnerability to the framework maintainers on January 4, 2024 including a minimal working example of the attack. Thus, while both protocols satisfy FEP-CPFA, neither can satisfy FEP-CCFA since they do not satisfy INT-CST.
@@
(lib)InterMAC. (lib)InterMAC is a protocol proposed by Boldyreva et al. (Boldyreva et al., 2012) and implemented by Albrecht et al. (Albrecht et al., 2019) which modifies the SSH Binary Packet Protocol in order to provide security properties associated with ciphertext lengths and fragmentation. The protocol transmits encrypted and authenticated fixed-sized chunks with the chunk size predetermined as a configuration parameter, padding to achieve this length if necessary. In response to our queries, the maintainer updated the libInterMAC repository with a complete implementation of the protocol on September 27, 2023, which is the version we tested. When a chunk fails to authenticate, libInterMAC terminates the connection in either direction as we verified in our experiments. We note that this protocol is not in use, and only encrypts the data transport phase of an SSH connection meaning that the existing InterMAC implementation is not properly a FEP, but we considered it valuable to include it here as an example of a distinct and reasonable approach to this problem from the others discussed here.
Obfs4. Obfs4 (obfs4, The obfourscator) is a fully encrypted transport protocol for Tor (Dingledine et al., 2004). It is a complex protocol that includes a key exchange and resists active probing attacks. The data transport phase of the protocol includes AEAD encrypted ciphertext blocks prepended by a two byte unauthenticated length field which is encrypted by a stream cipher. We installed Obfs4Proxy version 0.1.3 from the official Debian repository and used ptadapter (ptadapter Project Page., 2020) to run the transport as a standalone proxy. When a ciphertext fails to decrypt, the connection is immediately terminated by the receiver. This also occurs if an adversary modifies the length field in transit and the resulting expected ciphertext of an incorrect length fails to decrypt, which we were able to verify in our experiments. Thus while the protocol plausibly satisfies INT-CST (if a length field is modified, the protocol produces only channel closures/errors) it does not satisfy FEP-CCFA, since this close function is not secure (it requires knowledge of the encrypted lengths).
@@
OpenVPN-XOR. OpenVPN (OpenVPN Project., 2024) is a VPN protocol that can be patched to encrypt all outputs by XOR with the repetition of a pre-shared key (OpenVPN XORPatch., 2018). We analyzed OpenVPN version 2.6.8 as provided from the official Debian repository without the XOR patch applied (and assume that the XOR encryption should not modify message lengths) using the parameters set in the example configurations provided with the software, which deploy the proxy using the AES256-GCM cipher suite and with TLS mode turned on, with other configuration options mostly remaining as default. The obfuscation method in place is trivially identifiable and other work (Xue et al., 2022) has carefully studied its close behavior, which is to terminate the connection after an authentication failure. We explain the structure of OpenVPN messages as described in the official documentation (OpenVPN Network Protocol Documentation., 2024) under these default settings. OpenVPN messages are either data channel messages or control channel messages. Control channel messages are a minimum of 40 bytes if SHA-1 is used for the protocol’s HMAC (which is the default) and if the 4-byte timestamps intended to combat replay attacks are switched off (in our installation, they are on by default). This means adopting the default settings, control channel messages are 44 bytes. Version 2 data channel messages under these settings include a 26 byte header, including a two-byte length field. We did not observe Version 1 data channel messages in our experiments under the default configuration. OpenVPN typically encapsulates other protocols in the data channel, including their headers, but also directly sends messages of its own, including a fixed 16-byte ping message (explicitly defined in src/openvpn/ping.c Line 42), which was the smallest we observed in our experiments. Thus, the minimum length we identified in the TCP setting was 42.
@@
Obfuscated OpenSSH. Obfuscated OpenSSH (GitHub, 2010) is a patch to OpenSSH wherein key material is transmitted in the clear (or a PSK is used), after which the handshake protocol is fully encrypted (the binary packet protocol for SSH is fully encrypted without any modification). The approach of Obfuscated OpenSSH is used in the Psiphon (Psiphon., 2024) anti-censorship application, designed for Windows and mobile devices. We analyzed Obfuscated OpenSSH as downloaded from the official repository, which is a modification of OpenSSH 5.2_p1 but since the protocol includes complex handshake logic we ran it directly through a MiTM proxy to modify bytes in transit through the loopback interface and analyzed the resulting packet captures with Wireshark. OpenSSH terminates the connection immediately in both directions when a binary packet protocol message fails to authenticate, and we found in our experiments that this is no different after the obfuscation patch is applied. OpenSSH does not pad message lengths though this is permitted in the SSH specification. According to the RFC for the binary packet protocol, ”The minimum size of a packet is 16 (or the cipher block size, whichever is larger) bytes (plus ’mac’)” (Lonvick and Ylonen, 2006). For the initial messages, before the session key is negotiated, the MAC field is empty, meaning we expected the minimum to be 16 bytes. We ran the protocol in obfuscated and non-obfuscated configuration, and identified 16 byte messages appearing as part of the initial handshake in both circumstances. In the non-obfuscated captures, we identified this as a New Keys message, which indicates the switch to the recently negotiated session keys.
@@
kcptun. kcptun (kcptun, 2024) is a custom datastream transport protocol available in the official Debian repository, implemented via UDP datagrams and designed for very noisy networks. The protocol uses error correcting codes and sends many redundant messages to achieve a high probability that the sequence of messages can be reconstructed on the receiver side even in degraded network conditions. kcptun packets are by default encrypted using AES in Cipher Feedback mode, with a Cyclic Redundancy Check applied to the plaintext and prepended before encryption, providing non-cryptographic authentication, meaning that while the messages are fully encrypted, the protocol does not satisfy FEP-CCFA since it does not satisfy INT-CST. When the CRC is invalid, the packet is dropped, and kcptun measures the number of dropped packets and adjusts its internal parameters accordingly, significantly and predictably increasing the rate at which redundant messages are sent in direct proportion to the network packet loss. We wrote distinct software to proxy UDP messages, and did not notice a change in protocol behavior with increased induced packet loss beyond a predictable change in volume of packet transmissions. kcptun does not include padding and uses empty messages as keepalives which are the smallest that arose in our experiments, which are 16 (nonce) + 4 (CRC) + 24 (KCP headers) + 8 (Error Correction metadata) bytes for a total of 52 byte keepalive messages. The message format for kcptun messages is complex: a more thorough description of the headers, encapsulation format, and final message layout is given in the project page for the dependency kcp-go (KCP-Go project repository., 2024).
@@
SWGP-go. SWGP (Simple Wireguard Proxy., 2024) is a proxy designed for the Wireguard VPN protocol, intended to transform Wireguard into a datagram FEP. It behaves in two modes: one called “zero overhead”, where each Wireguard packet has its first 16 bytes encrypted via AES-ECB using a pre-shared key, and “paranoid”, which pads and AEAD encrypts each message. We assess the protocol in its paranoid mode. We deployed SWGP Version 1.5.0 configured with the default settings suggested by the maintainers on the project landing page. Since we suspected that running both multiple Wireguard and SWGP instances locally would be cumbersome, we ran our experiments through Docker-Compose (Docker Compose., 2024) and constructed three Docker images, a client image (including the client Wireguard instance and the client SWGP application), a router image which passed on messages transparently while recording them via TShark, and a server image. The client image also included a Python program to send an empty, raw IP packet through the interface every 5 seconds. Each experiment ran for 5 minutes and we assessed the resulting packet captures recorded by the router image. We observed in the source code that the paranoid protocol padded each message to a random length between one more than the packet’s minimum length and the MTU, which is set as a configuration option and is 1500 as default. Thus we hypothesized the minimum lengths to be the smallest messages in Wireguard, which are 32 byte keepalives, encapsulated with a 24 byte nonce and a 16 byte authentication tag, and a 2 byte length field (the encapsulation format is given in packet/paranoid.go, Line 18) with one required minimum padding byte for a total of 75, which is the smallest we observed in our experiments.